# Large Language Model-driven Meta-structure Discovery in Heterogeneous Information Network

### v1

@@ 3.5. Differential LLM Explainer
One prominent advantage of the LLM lies in its unparalleled ability for natural language generation. To harness this ability, we design a differential LLM explainer module to automatically generate human-comprehensible textual explanations elucidating the reasons behind the superior performance of discovered meta-structures. We design a prompting process in the chain-of-thought flavor (Wei et al., 2022b) to guide the differential LLM explainer in discerning the critical structural properties that contribute to performance enhancement. Specifically, the prompting process for analyzing a given meta-structure $T$ unfolds in the following two steps:
@@
Step 1: Structural Comprehension. We first sample a set of $n$ one-step neighbors for $T$, and translate each of them into a natural language sentence according to Method 3.1. Then, we prompt the differential LLM explainer to conduct a comprehensive analysis of both the original meta-structure and all its sampled neighbors. This involves breaking down each of them into meaningful sub-structures and identifying the sub-structure functions.
@@
Step 2: Performance Attribution. We first perform a quick evaluation of $T$ and each of the sampled neighbors separately by training a GNN with one structure at a time for downstream tasks. Then, we ask the differential LLM explainer to identify the presence/absence of beneficial/detrimental sub-structures in $T$. This involves a joint consideration of the evaluated performances and the structural analysis conducted in the previous step.
@@
Combination of these two steps empowers the differential LLM explainer to unravel the intricate connections between structural properties and performance, providing a comprehensive understanding of the discovered meta-structures. The effectiveness of this module is further justified by a pilot study involving human evaluators, which is elaborated in Section 4.4.2.
@@ 4.1.1. Datasets
We evaluate our proposed framework ReStruct on two important tasks in heterogeneous information network learning: recommendation and node classification.
For the recommendation task, the goal is to predict the existence of links between source nodes (e.g., users) and target nodes (e.g., businesses or movies). Under this task, we conduct experiments on three widely used real-world datasets: Amazon, Yelp, and Douban  [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding. Each dataset includes user ratings of items, which are converted to 0-1 binary labels with a threshold of $2$ to indicate the presence or absence of preference. Among the user-item pairs with the label ’1’, we randomly select half of them as positive pairs, which are further randomly split into train-validation-test sets with a ratio of 3:1:1. The other half is reserved for network construction so as to prevent label leakage. We take all user-item pairs with the label ’0’ as negative pairs, and also randomly split them into train-validation-test sets to pair each positive pair. If the total number of negative pairs is not enough, we randomly sample unconnected items until reaching the desired number. We adopt AUC (the area under the ROC curve) as the evaluation metric.
For the node classification task, the goal is to predict the correct labels of nodes of a given type, such as the genre of a movie. Under this task, we conduct experiments on two other widely adopted real-world datasets: ACM and IMDB. In these datasets, the classification targeted node type is papers (subjects) for ACM and movies (genres) for IMDB, respectively. We follow the data splits used in previous works (Ding et al., 2021; Li et al., 2023; Yun et al., 2019). We adopt the Macro-F1 score as the evaluation metric.
Details of all five datasets can be found in Appendix A.
@@ 4.1.2. Baselines
  • Hand-crafted meta-paths: HAN (Wang et al., 2019), which is a heterogeneous GNN that learns graph representation with multiple hand-crafted meta-paths and fuses them with a multi-head attention mechanism.
  • Automatically-searched meta-paths: RMSHRec (Ning et al., 2022), which adopts a reinforcement learning framework.
  • Automatically-searched meta-structures: (1) GEMS (Han et al., 2020), which searches meta-structures for recommendation on HINs with a genetic algorithm; (2) DiffMG (Ding et al., 2021), which formalized the problem into a neural architecture search flavor and search for meta-structures in a differentiable manner; (3) PMMM (Li et al., 2023), which further generalizes with multi-graph search.
@@ 4.1.3. Hyperparameter Settings
For our model, we run the algorithm for $30$ generations with a population size of $5$ and an elimination rate of $0.2$. When modifying each meta-structure, we randomly sample a set of $20$ candidates if there are too many of them from the one-step neighbors. When predicting meta-structure performances with the LLM predictor, we randomly sample a set of $30$ evaluation records from the performance pool to enable a few-shot learning paradigm through prompting with examples. For the LLM, we use the GPT-4 model by calling the OpenAI API. For all baseline models, we follow the implementation released by the authors. We fix the hidden dimensions to $64$ for all evaluated models, and tune other hyperparameters including learning rate, weight decay, and dropout by referring to the performances on the validation set. To reduce the noise brought by randomness during program execution, for each combination of (model, task, dataset), we run experiments with $10$ different random seeds and report the average performance with standard deviation.
@@ 4.2. Comparison on Recommendation
In Table 1, we report the experimental results of our proposed framework on the recommendation task compared to baselines. First, we observe that models using meta-paths (HAN, RMSHRec) exhibit substantially poorer performances than those using meta-structures (all the rest), confirming that the stronger expression capability of meta-structures is desired for heterogeneous graph learning. Second, ReStruct consistently achieves the best performance across three datasets, showcasing the effectiveness of our framework in identifying meaningful and useful structures in various HINs. Specifically, the performance gain over GEMS confirms that LLM-guided “targeted mutation” converges to better solutions that pure random mutation in a classic genetic framework.
HAN RMSHRec GEMS DiffMG PMMM ReStruct
Amazon 63.53$\pm$0.24 61.80$\pm$0.15 63.76$\pm$0.47 73.27$\pm$0.20 73.78$\pm$0.04 75.27$\pm$0.19
Yelp 74.54$\pm$0.09 59.39$\pm$0.81 75.46$\pm$0.28 77.63$\pm$0.40 76.76$\pm$0.17 84.04$\pm$0.24
Douban 85.12$\pm$0.12 79.72$\pm$0.32 90.84$\pm$0.10 93.94$\pm$0.04 94.31$\pm$0.02 94.49$\pm$0.03
Table 1. AUC (%) of recommendation on three datasets.
@@ 4.3. Comparison on Node Classification
In Table 2, we report the experimental results of our proposed framework on the node classification task compared to baselines. First, HAN and RMSHRec do not achieve desirable performances. Specifically, HAN is limited by its heavy reliance on hand-crafted meta-structures, while RMSHRec suffers from high instability presumably due to the limitation accompanied by the reinforcement learning framework. Second, DiffMG and PMMM demonstrate improved performances, showcasing the advantage brought by meta-structures over meta-paths, as well as the NAS searching framework. Finally, the meta-structure discovered by ReStruct achieves the best performance on both ACM and IMDB datasets.
HAN RMSHRec DiffMG PMMM ReStruct
ACM 91.20$\pm$0.25 90.32$\pm$0.74 92.65$\pm$0.15 92.76$\pm$0.14 92.82$\pm$0.23
IMDB 55.09$\pm$0.67 56.72$\pm$2.60 61.04$\pm$0.56 61.69$\pm$0.40 63.32$\pm$0.62
Table 2. Macro F1 scores (%) of node classification on two datasets.
4.4. Explainability Analysis of Discovered Meta-Structures
@@ 4.4.1. Visualization of Discovered Structures and LLM-generated Explanation
In Figure 5, we present $3$ discovered meta-structures that are among the top-performing ones on the Yelp dataset for the recommendation task, along with a summary text explaining the structural attributes underlying their outstanding performances, which are essentially output from our differential LLM explainer further condensed for clarity. We highlight the LLM-identified good (relevant) sub-structures in red, while the bad (distracting) sub-structures in blue. This visualization facilitates a comprehensive understanding of each structure’s composition. For example, the meta-structure in Figure 5 (a) contains critical yet simple sub-structures describing the geographical, social, and business category contexts of user behavior. While more nodes and edges can create complex relationships, they are not always desirable for HIN learning, as showcased by the identified distracting sub-structures. These structures, though challenging to handcraft, carry semantically meaningful explanations that remain accessible with textual assistance.
4.4.2. Pilot Study of Human Evaluation
We conduct a pilot study with $73$ human participants, more specifically, graduate students with domain knowledge of HIN research, to evaluate the level of explainability provided by our framework compared to baselines. We structure the study around two sets of questions. The first set aims to compare the explainability of the best-performing meta-structure discovered by our model with those found by several baseline models, including GEMS, DiffMG, and PMMM. Participants are asked to select one meta-structure from a set of meta-structures, each being the best one found by a specific model, as documented in the original papers (Han et al., 2020; Ding et al., 2021; Li et al., 2023). The second set of questions investigates the comprehension gain brought by the textual explanation generated by our differential LLM explainer, as described in Method 3.5, coined the Differential Explanation. As a baseline, we include the textual explanation generated by directly prompting an LLM to explain the reasons behind a given meta-structure’s strong performance, without undergoing the two-step prompting process, coined the Non-Differential Explanation. For three meta-structures among the best ones discovered by our model on the Yelp HIN, participants are presented with both types of explanations. They are then asked to determine which explanation is more helpful in enhancing their (1) understanding of how this meta-structure is constructed, and (2) insights on how to design a better one. The complete questionnaire utilized in our study is available in Appendix B.
@@ 4.4.2. Pilot Study of Human Evaluation
4.5. Ablation Study
To assess the efficacy of the designed grammar translator and LLM decision modules, we conduct ablation studies involving two model variations: (1) No Grammar: In this scenario, the grammar translator module is excluded from the framework. LLM agents are prompted to predict candidate performances and select the best candidates directly based on the matrix expression of meta-structures (Han et al., 2020). (2) No Grammar, No LLM Decision: This variation involves the removal of both the grammar translator module and the LLM decision modules from the framework. Candidates are randomly selected to advance to the next generation.
As shown in Figure 8, both ablation models exhibit inferior performance compared to the full model. Specifically, among the three models, No Grammar, No LLM Decision is more prone to getting stuck in local sub-optimal solutions. The LLM selection module facilitates a faster escape from the sub-optimal, showcasing the advantage of knowledge-guided exploration. Ultimately, the full model navigates the search space with significant strides, suggesting that our grammar encoding of the meta-structures contributes to performance enhancement by enabling LLM agents to better grasp the semantic meanings of complex meta-structures. In summary, these results affirm that both designed components contribute considerably to a better meta-structure discovery process.
Figure 8. Comparison of best-performing meta-structure found by each generation.
@@ 5. Related Works
5.1. Identifying Meta-structure in HIN
@@
In the past decade, heterogeneous information networks (HIN) gradually gained popularity for their ability to capture the complex relations between multi-typed nodes, which play important roles in various research areas such as information retrieval and social network modelling (Shi et al., 2016). Meta-path, a pre-defined path template of relation sequences, was proposed to measure the similarities between nodes on HIN (Sun et al., 2011). It allows search algorithms like PathSim (Sun et al., 2011) to find peer nodes that are connected by paths with different semantic meanings. The concept of meta-path was later extended to a more general form of meta-structure (Huang et al., 2016), where the relation patterns between connected nodes can be characterized as a directed acyclic graph. Previous works have found meta-structures useful for boosting machine learning performance on HINs (Jiang et al., 2017). However, early works were based on carefully hand-crafted meta-structures, which heavily relied on domain knowledge. To address this challenge, several more recent works proposed to automate meta-structure design with heuristic algorithms (Meng et al., 2015), reinforcement learning (Yang et al., 2019), evolutionary search (Han et al., 2020) and differentiable structure learning (Ding et al., 2021). Different from previous efforts of automatic meta-structure design, we are first to leverage the emergent reasoning ability of LLMs (Wei et al., 2022a). We design novel LLM modules for the automatic generation, evaluation, and explanation of novel meta-structures, which are proven effective in several representative datasets.
5.2. Deep Learning on HIN

### v2

@@ 3.5. Differential LLM Explainer Agent
One prominent advantage of the LLM lies in its unparalleled ability for natural language generation. To harness this ability, we design a differential LLM explainer agent (abbreviated as “explainer” below) to automatically generate human-comprehensible textual explanations that elucidate the reasons behind the superior performance of discovered meta-structures. To guide the explainer in discerning the critical structural properties that contribute to performance enhancement, we design a prompting process in the chain-of-thought flavor (Wei et al., 2022b) . Specifically, for analyzing a given meta-structure $T$, it unfolds in the following two steps:
@@
Step 1: Structural Comprehension. We begin by sampling a set of $n$ one-step neighbors for the meta-structure $T$, and translate each of them into a natural language sentence according to Method 3.1. Then, we prompt the differential LLM explainer to conduct a comprehensive analysis of both $T$ and all its sampled neighbors. This process involves breaking down each of them into meaningful sub-structures and identifying the functions os these sub-structures.
@@
Step 2: Performance Attribution. We first perform a quick evaluation of $T$ and each of the sampled neighbors separately by training a GNN with one structure at a time for downstream tasks. Then, we ask the differential LLM explainer to identify the presence/absence of beneficial/detrimental sub-structures in the meta-structure $T$. This attribution process involves a joint consideration of the evaluated performances and the structural analysis conducted in the previous step.
@@
The combination of these two steps empowers the explainer to unravel the intricate connections between structural properties and empirical performance, providing a comprehensive understanding of the discovered meta-structures. The effectiveness of this module is further justified by a user study involving human evaluators, which is elaborated in Section 4.4.2.
@@ 4.1.1. Datasets
We evaluate ReStruct on two important tasks in HIN learning: recommendation and node classification, each with four datasets covering different fields. Detailed statistics of all eight datasets can be found in Appendix A.
In the recommendation task, our goal is to predict the existence of links between source nodes (e.g., users) and target nodes (e.g., items or businesses). We conduct experiments on four widely-used real-world datasets: Amazon, Yelp, Douban Movie (abbreviated as ”Douban”), and LastFM [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding. For datasets including user ratings of items, the ratings are converted to 0-1 binary labels according to a threshold of $2$. A label of 1 indicates the presence of preference, while 0 indicates the absence. Among the user-item pairs with the label ’1’, we randomly select half of them as positive pairs, which are further randomly split into training-validation-testing sets with a ratio of 3:1:1. The other half is reserved for network construction so as to prevent label leakage. We take all user-item pairs with the label ’0’ as negative pairs, and also randomly split them into train-validation-test sets to pair each positive pair. If the number of negative pairs is insufficient, we randomly sample unconnected items until reaching the desired number. The evaluation metric used in these experiments is AUC (Area Under the ROC Curve), which measures the model’s ability to rank positive instances higher than negative instances.
In the node classification task, our goal is to predict the labels of nodes belonging to a specific type, such as determining the genre of a movie. To evaluate the effectiveness of our approach, we perform experiments on four widely-adopted real-world datasets: ACM, IMDB, DBLP, and OAG-NN. In these datasets, the classification targets correspond to the subjects of papers in ACM, the genres of movies in IMDB, the research areas of authors in DBLP, and the published venues of papers in OAG-NN, respectively. For ACM, IMDB, and DBLP, we follow the data splits used in previous works (Ding et al., 2021; Li et al., 2023; Yun et al., 2019). For OAG-NN (Hu et al., 2020), we filter the published venues with more than 100 recorded papers, and randomly split the dataset into training-validation-testing sets by 3:1:1. The evaluation metric used in these experiments is the Macro-F1 score, which measures the performance of the classification model in terms of precision and recall.
@@ 4.1.2. Baselines
  • Hand-crafted meta-paths: (1) metapath2vec (Dong et al., 2017), which trains a skip-gram model with meta-path guided random walks; (2) HIN2Vec (Fu et al., 2017), which learns latent vectors by jointly training for multiple prediction tasks; (3) HAN (Wang et al., 2019), which is a heterogeneous GNN that learns graph representation with multiple hand-crafted meta-paths and fuses them with a multi-head attention mechanism; (4) HERec (Shi et al., 2018), which combines random walks with an extended matrix factorization model.
  • Automatically-searched meta-paths: RMSHRec (Ning et al., 2022), which adopts a reinforcement learning framework to search for meta-paths.
  • Automatically-searched meta-structures: (1) GEMS (Han et al., 2020), which employs a genetic algorithm; (2) DiffMG (Ding et al., 2021), which adopts a neural architecture search manner and searches for meta-structures in a differentiable manner; (3) PMMM (Li et al., 2023), which further generalizes DiffMG with multi-graph search.
@@ 4.1.3. Hyperparameter Settings
For our model, we run the algorithm for $30$ generations with a population size of $5$ and an elimination rate of $0.2$. When modifying each meta-structure, we randomly sample a set of $20$ candidates if there are too many of them from the one-step neighbors. When predicting meta-structure performances with the few-shot LLM predictor, we randomly sample $30$ records from the performance pool to fuel the few-shot learning paradigm. To implement the LLM agents, we use the GPT-4 model by calling the OpenAI API [^2] https://platform.openai.com/docs/models/, while robustness analysis with other LLM models are also carried out (see Section 4.5). We employ the DGL implementation of HAN [^3] https://github.com/dmlc/dgl/tree/master/examples/pytorch/han. For all the other baseline models, we follow the implementation released by the authors. We fix the hidden dimensions to $64$ for all evaluated models, and tune other hyperparameters including learning rate, weight decay, and dropout by referring to the performances on the validation set. To reduce the noise brought by randomness during program execution, for each combination of (model, task, dataset), we run experiments with $10$ different random seeds and report the average performance with standard deviation.
@@ 4.2. Comparison on Recommendation
In Table 1, we report the experimental results of ReStruct on the recommendation task compared to baselines. First, we observe that models using meta-paths generally exhibit substantially lower performances than those using meta-structures, confirming that the stronger expression capability of meta-structures is desired for heterogeneous graph learning. Second, ReStruct consistently achieves the best performance across four datasets, showcasing the effectiveness of our framework in identifying meaningful and useful structures in various HINs. In particular, the performance gain over GEMS confirms that the LLM-guided “targeted mutation” converges to better solutions than pure random mutation in a classic genetic framework.
Table 1. AUC (%) of recommendation on four datasets.
metapath2vec HIN2Vec HAN RMSHRec HERec GEMS DiffMG PMMM ReStruct
Amazon 56.88$\pm$0.27 58.66$\pm$0.12 60.03$\pm$0.48 61.80$\pm$0.15 70.55$\pm$0.05 63.76$\pm$0.47 73.27$\pm$0.20 73.78$\pm$0.04 75.27$\pm$0.19
Yelp 52.29$\pm$0.59 68.06$\pm$2.31 63.88$\pm$2.32 59.39$\pm$0.81 68.37$\pm$0.34 75.46$\pm$0.28 77.63$\pm$0.40 76.76$\pm$0.17 84.04$\pm$0.24
Douban 52.84$\pm$0.00 85.95$\pm$0.06 63.27$\pm$0.38 79.72$\pm$0.32 92.22$\pm$0.00 90.84$\pm$0.10 93.94$\pm$0.04 94.31$\pm$0.02 94.49$\pm$0.03
LastFM 68.57$\pm$0.10 69.62$\pm$2.00 72.93$\pm$2.42 81.94$\pm$0.22 79.64$\pm$0.06 78.23$\pm$0.08 82.53$\pm$0.11 82.88$\pm$0.07 85.21$\pm$0.09
@@
[…]
Discovered meta-structures on Yelp and the corresponding generated natural language explanations.
@@ 4.3. Comparison on Node Classification
In Table 2, we report the experimental results of ReStruct on the node classification task compared to baselines. First, metapath2vec, HIN2Vec, and HAN do not achieve desirable performances, mainly due to their heavy reliance on hand-crafted meta-structures. Second, DiffMG and PMMM demonstrate improved performances, showcasing the advantage brought by meta-structures over meta-paths, as well as the NAS searching framework. Finally, ReStruct achieves the best performance on ACM, IMDB, and OAG datasets. It closely aligns with state-of-the-art results on the DBLP dataset, where Macro F1 scores already exceed $94\%$ – a level challenging to significantly surpass.
Table 2. Macro F1 scores (%) of node classification on four datasets.
metapath2vec HIN2Vec HAN DiffMG PMMM ReStruct
ACM 67.13$\pm$0.50 80.75$\pm$0.77 91.20$\pm$0.25 92.65$\pm$0.15 92.76$\pm$0.14 92.82$\pm$0.23
IMDB 40.82$\pm$1.48 48.16$\pm$0.44 55.09$\pm$0.67 61.04$\pm$0.56 61.69$\pm$0.40 63.32$\pm$0.62
DBLP 89.93$\pm$0.45 90.58$\pm$0.62 92.13$\pm$0.26 94.45$\pm$0.15 94.69$\pm$0.10 94.09$\pm$0.36
OAG-NN 27.61$\pm$1.65 47.13$\pm$1.14 45.80$\pm$5.84 37.67$\pm$3.13 30.18$\pm$3.87 47.52$\pm$1.56
4.4. Explainability Analysis
@@ 4.4.1. Visualization of Discovered Structures and LLM-generated Explanation
Figure 5 showcases $3$ discovered meta-structures that are among the top-performing ones on the Yelp dataset for recommendation, each with a summary text explaining the structural attributes underlying their outstanding performances. These summaries are generated by our differential LLM explainer and further condensed for clarity. We highlight the LLM-identified good (relevant) sub-structures in red, and the bad (distracting) sub-structures in blue. This visualization facilitates a comprehensive understanding of each structure’s composition. For example, the meta-structure in Figure 5 (a) contains critical yet simple sub-structures describing the geographical, social, and business category contexts of user behavior. While more nodes and edges can create complex relationships, they are not always desirable for HIN learning, as showcased by the identified distracting sub-structures. These structures, though challenging to handcraft, carry semantically meaningful explanations that remain accessible with textual assistance.
4.4.2. User Study for Human Evaluation
We conduct a user study to evaluate the explainability provided by our framework compared to baselines from a human perspective. As our framework targets HIN researchers and engineers as potential users, we recruit $73$ graduate students with domain knowledge of HIN research as our participants. To further ensure participants’ solid understanding of the survey’s processes and questions, we provide clear explanations of key concepts such as HIN, meta-path and meta-structure, supported by illustrative examples at the beginning of the survey. We structure the study around two sets of questions. The first set of questions is designed to assess the inherent comprehensibility of generated meta-structures without textual explanations. To this end, we present the visualizations of the best meta-structures discovered by our model alongside those from three baseline models (GEMS, DiffMG, and PMMM) (Han et al., 2020; Ding et al., 2021; Li et al., 2023), and ask the participants to select the most comprehensible one from their point of view. The second set of questions is designed to assess the comprehension gain brought by the textual explanation generated by our differential LLM explainer (see Method 3.5), coined the Differential Explanation. As a baseline, we include a Non-Differential Explanation, generated by directly prompting an LLM to explain the reasons behind a meta-structure’s strong performance without undergoing the two-step prompting process. For three meta-structures discovered by our model on the Yelp HIN, participants are presented with both types of explanations. They are then asked to determine which one is more helpful in enhancing their understanding of how the meta-structure is constructed and gaining insights on how to design a better one. By engaging participants in head-to-head comparisons, we aim to gather valuable feedback on the relative helpfulness of each explanation type. Before starting the survey, we also carried out a pilot study (Rattray and Jones, 2007) with $5$ participants to ensure the clarity of the questions and visualizations, and none of them expressed confusion or difficulty in understanding these materials. The complete questionnaire utilized in our study is available in our GitHub repository.
@@ 4.4.2. User Study for Human Evaluation
[…]
Human evaluated explainability of the best meta-structures discovered by different models.
@@
[…]
Explainability gain brought by LLM-generated differential meta-structure explanation compared to non-differential explanation.
4.5. Robustness Analysis
To test the robustness of our method, we first analyze its performance variation under prompt perturbation. Following previous studies (Sun et al., 2023a; Moraffah et al., 2024; Zhang et al., 2024), we ask ChatGPT to paraphrase our prompts while maintaining the key module designs in the framework such as grammar translation and LLM-guided performance prediction, and use them to replace the original prompts in the ReStruct framework. As shown in Table 3, we find that prompt paraphrasing does not have a clear impact on model performance, as long as the key designs are kept, the importance of which has been validated in the previous section.
Table 3. Model robustness to prompt perturbation with ChatGPT (taking recommendation on Yelp as an example).
Index Prompt for Paraphrasing AUC (%)
0 (Original) - 84.04$\pm$0.24
[…]
Second, we analyze ReStruct’s robustness by changing the underlying LLM. Specifically, we conduct experiments with nine different LLMs from three popular series (GPT, Mistral, and GLM), with different training data, training methods, and parameter sizes. As shown in Table 4, the model performances across these LLMs consistently and significantly outperform all baselines, demonstrating the robustness of our framework against LLM versions.
Table 4. Model performance with different LLM versions (taking recommendation on Yelp as an example).
Model AUC (%)
gpt-4-1106-preview (original) 84.04$\pm$0.24
gpt-3.5-turbo-1106 84.03$\pm$0.11
mistral-tiny-2312 83.65$\pm$0.17
mistral-small-2312 83.89$\pm$0.09
mistral-small-2402 83.96$\pm$0.12
mistral-medium-2312 84.04$\pm$0.16
mistral-large-2402 83.87$\pm$0.13
glm-3-turbo 84.12$\pm$0.06
glm-4 83.76$\pm$0.07
@@ 5. Related Works
5.1. Identifying Meta-structures on HINs
@@
Over the past decade, HINs have gained popularity for their ability to capture the complex relations between multi-typed nodes, which play important roles in various research areas such as information retrieval and social network modelling (Shi et al., 2016). Meta-path, a pre-defined path template of relation sequences, was proposed to measure the similarities between nodes on HIN (Sun et al., 2011). It allows search algorithms like PathSim (Sun et al., 2011) to find peer nodes that are connected by paths with different semantic meanings. The concept of meta-path was later extended beyond the linear relationship to a more general form of meta-structure (Huang et al., 2016), where the relation patterns between connected nodes can be characterized as a directed acyclic graph. Previous works have found meta-structures useful for boosting machine learning performance on HINs (Jiang et al., 2017). However, early works were based on carefully hand-crafted meta-structures, which heavily relied on experts’ domain knowledge. To address this challenge, several recent works proposed to automate meta-structure design with heuristic algorithms (Meng et al., 2015), reinforcement learning (Yang et al., 2019), evolutionary search (Han et al., 2020) and differentiable structure learning (Ding et al., 2021). Different from previous efforts of automatic meta-structure design, we are first to leverage the emergent reasoning ability of LLMs (Wei et al., 2022a) for this task. We design novel LLM agents for the automatic generation, evaluation, and explanation of novel meta-structures, which are proven effective in eight representative datasets.
5.2. Deep Learning on HINs